An airplane flying against the wind went 1500 miles in 5 hours and returned with the wind in 3 hours.Assume the wind rate didnt change,find the rate of the airplane and the wind.
Answer by mananth(16946)

Plane speed =x mph
wind speed =y mph
against wind 5 hours
with wind 3 hours

Distance against 1500 miles distance with 1500 miles
t=d/r against wind -
1500 / ( x - y )= 5
5 ( x - y ) = 1500
5 x - 5 y = 1500 ....................1

1500 / ( x + y )= 3
3 ( x + y ) = 1500
3 x + 3 y = 1500 ...............2
Multiply (1) by 1
Multiply (2) by 1
we get
5 x + -5 y = 1500
3 x + 3 y = 1500
8 x = 3000
/ 8
x = 375.00 mph

plug value of x in (1) y
5 x -5 y = 1500
1875 -5 -1875 = 1500
-5 y = 1500
-5 y = -375 mph
y = 75.00
Plane speed 375.00 mph
wind speed 75.00 mph

m.ananth@hotmail.ca